Alt Ctri 8 =om 2) Find F = Vi0 Oof the following cash flow 100 100 200 300 400 500 600 0 800 900 1000 1 2 3 4 5 6 7 |10 8 l060 a) 996-1001 b) 1001 - 1006 6480-6540 c) 1006-1011 6 540-6600 d) 1011 10166600- 6660 e) None of the above 600
Thu 2) Find F = V10 of the following cash flow Vio is the futwe value in yean lo 16 100 100 200 300 400 500 600 0 800 900 1000 1 2 3 5 6 7 8 10 l060 6420-6480 a) 996-1001 b) 1001 1006 6480-6540 c) 1006-1011 6540-6600 d) 1011 1016 6600- 6660 e) None of the above 600 De
i8% 2) Find F= Vio of the following cash flow (Vio is the futue value in yean lo 1000 800 900 0 600 500 100 100 200 400 300 10 7 8 2 3 6 5 l060 a) 996-1001 420-480 b) 1001-1006 6480-6540 c) 1006-1011 6540-6600 d) 1011 -10166600-6660 e) None of the above 100 (PIG,48 Do
Answer #1

As question is not so clear.I assume that you are asking me to find out what is the future value or value of the cash flows at the end of 10th year..

Assuming cashflows in between are yearly cashflows you are investing every cashflow at 8 % interest rate .. value of the cashflows matured along with interest is as follows:

Cashflow received at year 0 - is invested for 10 years will be matured at the end of 10th year = 100(1.08) ^10= 1080

Cashflow at end of year 1 will be invested for 9 years= 100(1.08)^9=972

End of year 2 = 200(1.08)^8= 1728

Y3 = 300(1.08)^7=2268

Y4= 400(1.08)^6= 2592

Y5 = 500(1.08)^5=2700

Y6=600(1.08)^4=2592

Y7= 0

Y8=800(1.08)^2=1728

Y9= 900(1.08)^1= 972

Y10= 1000

By adding all fv at end of y10 = 17632